OpenStudy (anonymous):

WILL FAN AND MEDAL A group of roommates decide to split the cost of their groceries. If g is the amount spent on groceries and n is the number of roommates, the formula c= g/n determines each person’s share of the cost. If each person’s share was $104.50 and 6 roommates live in the house, what was the cost of the groceries?

OpenStudy (thecurious):

104.50*6 thats the cost of groceries!

OpenStudy (texaschic101):

6 * 104.50

OpenStudy (sidsiddhartha):

share of each person=g/n 104.50=g/n g=104.5*6=627